Tiara is driving from her house to her grandmother's house that is 178 miles away. She is driving at rate of 65 miles per hour. If she wants to know how long the trip will take she should.....
Answer by mukhopadhyay(490)

Use formula Distance(D) = Rate(R)*Time(T)
D = RT
=> 178 = 65t
=> t = 178/65 Hours = 2 48/65 hours